# Boltzmann priors for Implicit Transfer Operators

### v1

@@ A.4.1 Boltzmann Priors for dataset generation
For different number of trajectories, $n$, we train $5000/n$ ITO models on $n$ Prinz potential trajectories of length 150. Trajectories do not overlap among different trainings. The maximum model lag is 100 steps.
@@
A.4.2 BoPITO samples efficiently long-term dynamics in data-sparse scenarios
@@
A.4.3 Interpolating between models trained on off equilibrium data and the Boltzmann distribution
@@
A.4.4 Architectural and training details
@@
Table 1: Architectural and training parameters of MB-ITO models (a) and SE3-ITO (b).
A.5 Prinz potential
@@
Figure 8: Implied time-scales (a) and dynamic observable under the identity function (b) for the Prinz Potential. Errors in (b) are smaller than the dots.
A.6 Alanine Dipeptide
@@
Figure 9: Implied time-scales for Alanine Dipeptide computed with a Markov State Model on torsional angles $\phi$ and $\psi$ (a) and dynamic observable in eq. 22 (b). Errors in (b) are smaller than the dots.
@@
A.7 Metrics
@@
A.8 Energies for Alanine Dipeptide
@@
A.9 Fitting $\hat{\lambda}$
The hyperparameter $\hat{\lambda}$ controls the time-scale at which the BoPITO model transitions to sampling the equilibrium model. If $\hat{\lambda}$ is too small, the model may prematurely relax to equilibrium, limiting its ability to capture non-equilibrium dynamics. Conversely, if $\hat{\lambda}$ is too large, the benefits of the BoPITO framework could diminish. Therefore, careful tuning of $\hat{\lambda}$ can be crucial for optimal performance.
As illustrated in Figure A.9, grid-search hyperparameter tuning can select an appropriate $\hat{\lambda}$ for a single-system model. Too small values of $\hat{\lambda}$ can hinder the model’s ability to capture long-time-scale dynamics, leading to increased loss. We recommend choosing the smallest $\hat{\lambda}$ that yields a plateau in the $\hat{\lambda}$-loss curve (elbow rule), as demonstrated in Figure A.9 (b). However, practitioners should be cautious about sampling longer lags than the implied time-scale defined by $\hat{\lambda}$ if they cannot guarantee the system relaxes to equilibrium for those time-scales.
Figure 12: Average validation loss against training epochs for the Prinz potential for different values of $\hat{\lambda}$ (a) and average final loss versus $\hat{\lambda}$ (b). Averages are taken w.r.t. 10 runs. Final losses are computed as the average between epochs $190$ and $200$. The greatest eigenvalue of the system under our simulation parameters is $0.994$.
A.10 Correlation function normalization

### v2

@@ A.9.1 Boltzmann Priors for dataset generation
For different numbers of trajectories, $n$, we train $5000/n$ ITO models on $n$ Prinz potential trajectories of length 150. Trajectories do not overlap among different trainings. The maximum model lag is 100 steps.
@@
A.9.2 BoPITO samples efficiently long-term dynamics in data-sparse scenarios
@@
Chignolin
[…]
A.9.3 Interpolating between models trained on off equilibrium data and the Boltzmann distribution
@@
A.9.4 Architectural, training and inference details
@@
Inference ODE steps $100$ /$50$
Table 1: Architectural and training parameters of MB-ITO models (a) and SE3-ITO (b). Batch sizes and inference ODE steps refer to Alanine Dipeptide and Chignolin experiments respectively.
A.10 Prinz potential
@@
Figure 10: Implied time-scales (a) and dynamic observable under the identity function (b) for the Prinz Potential. Errors in (b) are smaller than the dots. In (a), the implied time-scales of the $5$ slowest processes in the system are computed for different lags. The color order, from longest to shortest implied time-scale, is blue, orange, green, red, and purple.
A.11 Alanine Dipeptide
@@
Figure 11: Implied time-scales for Alanine Dipeptide computed with a Markov State Model on torsional angles $\phi$ and $\psi$ (a) and dynamic observable in eq. 26 (b). In (a), the implied time-scales of the $5$ slowest processes in the system are computed for different lags. The color order, from longest to shortest implied time-scale, is blue, orange, green, red, and purple.
@@
A.12 Chignolin
Chignolin (cln025) is a fast folding protein with $10$ residues, $166$ atoms and $93$ heavy atoms. We use molecular dynamics data previously reported by Lindorff-Larsen et al. (2011). The data is proprietary but available upon request for research purposes. The simulations were performed in explicit solvent with a $2.5$ fs time-step and the positions was saved at $200$ ps intervals. We extract all heavy atoms positions from the simulations and train models on this data. We use the reaction coordinate proposed in Lindorff-Larsen et al. (2011) as to define a dynamic observable,
$a(\mathbf{x})=b(\mathbf{x})=\frac{\sum_{i=1}^{N_{\text{res}}}\sum_{j>i}^{N_{i}}\frac{1}{1+e^{10(d_{ij}(\mathbf{x})-d_{ij}^{*}-1)}}}{\sum_{i=1}^{N_{\text{res}}}N_{i}},$ (27)
where the first sum in the numerator iterates over all $N_{\text{res}}$ residues in the protein, while the second sum considers the $N_{i}$ native contacts of residue $i$ separated by at least seven residues in the primary sequence. We define native contacts as residue pairs separated by at least seven residues in the primary sequence and with $C_{\alpha}$ atoms closer than 10 Å in the native structure. Moreover, $d_{ij}(\mathbf{x})$ and $d_{ij}^{*}$ represent the distances between the $C_{\alpha}$ atoms of residues $i$ and $j$ in the structure $\mathbf{x}$ and the native structure, respectively. This observable quantifies the protein’s foldedness, with values ranging from 0 (unfolded) to 1 (folded). In Figure 13 (a) we show the evolution of eq. 27 on a subset of the reference simulation data and we observe how the protein undergoes transformations between the folded and unfolded states. In Figure 13 (b) we visualize the corresponding dynamic observable.
Figure 13: Time evoluation of ‘foldedness’ in a subset of the reference MD simulation (a) and dynamic observable in eq. 27 (b).
A.13 Metrics
@@
A.14 Energies for Alanine Dipeptide
@@
A.15 Correlation function normalization